The Research Methods and Statistic module is almost over. Assume that the total marks were at an average of μ= 18 per assessment, and that the distribution of total marks are normally distributed with σ= 10. What proportion of students would have marks between 10 and 24? 


1
Expert's answer
2021-06-07T12:13:00-0400

Let "X=" the mark of student: "X\sim N(\mu, \sigma^2)"

Given "\mu=18, \sigma=10."

"P(10<X<24)=P(X<24)-P(X\leq 10)"

"=P(Z<\dfrac{24-18}{10})-P(Z\leq \dfrac{10-18}{10})"

"=P(Z<0.6)-P(Z\leq -0.8)"

"\approx0.7881446-0.2742531\approx0.5139"